### v1

Abstract
This is a critical discussion of theoretical and descriptive deficiencies of the recently used ”flat Odderon” model.
A.A. Logunov Institute for High Energy Physics
NRC ”Kurchatov Institute”, Protvino, RF
Introduction
For more than half a century, the properties and significance of Odderon, the C-odd partner of Pomeron and presumably asymptotically subleading w.r.t. the latter, have been discussed with varying intensity. By now, one can find a whole spectrum of opinions regarding the Odderon.
Experiment:
  • even if it exists, its influence at high energies is insignificant;
  • the Odderon effect has already been discovered at the LHC in the region of small momentum transfers;
  • Odderon is unnoticeable at small transfers, but is clearly expressed in the region of the dip ($pp$) and shoulder ($\bar{p}p$);
  • It must be sought in processes other than $pp$ and $\bar{p}p$ elastic scattering.
Theory:
  • Odderon is the position of the rightmost singularity of the C-odd amplitude in the j-plane;
  • Odderon is a Regge pole.
  • Odderon is a Regge cut.
For a brief overview of the current state of the Odderon problem, see, for example, [1].
In our note we are more specific and would like to touch upon one extreme case of realization of the Odderon as a ”flat Regge trajectory” [2] (actually a fixed pole) with help of which several ”two-channel” eikonal model variants were tested with a data description on $pp$-scattering at $\sqrt{s}=50\div 13000$ GeV and for $-t<0.1\;GeV^{2}$. With use of some special prescription for calculation of the $\chi^{2}$ the authors of the paper in question claim quite a reasonable description of the data with $\chi^{2}/ndf=1.11$ and indicate that the absence of the Odderon contribution worsens the quality of the description though this contribution is small. The sign of the Odderon was also defined and the value of the Odderon coupling $\beta_{\mathcal{O}}(t)$ at $t=0$ is estimated to be less than that of the Pomeron. Moreover, the contribution of the Odderon to the ratio $\Re T^{pp}(s,0)/\Im T^{pp}(s,0)$ was found to be an order of magnitude smaller than that proclaimed by the TOTEM collaboration.
All these problems are of significant interest and importance, which, in fact, prompted this note.
In relation to theory we will limit ourselves to discussing only the Odderon part, while the evaluation of the statistical part applies to the entire work as a whole.
Some preliminaries
Keeping in mind the accompanying ”unitarization” (eikonal representation) and for reference we remind the standard expression for the Odderon as a ”Born amplitude” $\mathcal{O}(s,t)$  [^3] For the sake of certainty, we will assume that we are talking about proton-proton scattering. We do not touch here on the question of the sign of the Odderon contribution, which is irrelevant for the subject of our note. defined by the negative signature Regge pole at $J=\alpha(t)$ (see e.g. Ref. [3])
$\mathcal{O}(s,t)=\pi\frac{\alpha^{{}^{\prime}}(t)[2\alpha(t)+1]}{2\sin\left[\pi\alpha(t)\right]}\left[1-e^{-i\pi\alpha(t)}\right]\Gamma_{\alpha(t)}^{2}(t)P_{\alpha(t)}(-z_{t})=$ (1)
$=\pi\alpha^{{}^{\prime}}(t)\left[\alpha(t)+\frac{1}{2}\right]\left\{i+\mbox{tg}\left[\frac{\pi\alpha(t)}{2}\right]\right\}\Gamma_{\alpha(t)}^{2}(t)P_{\alpha(t)}(-z_{t})$
where $\Gamma_{\alpha(t)}(t)$ is the spin $J$ meson vertex $\Gamma_{J}(t)$ at $J=\alpha(t)$ and $z_{t}=1-2s/(4m_{J}^{2}-t)$. When moving to the $t$-channel, $\Re\alpha(t)\rightarrow J=1,3,...$, and the propagator of the C-odd meson of (generally complex) mass $m_{J}$ is reproduced:
$\mathcal{O}(s,t)\sim\frac{\Gamma_{J}^{2}(t)}{m^{2}_{J}-t}P_{J}(z_{t}).$ (2)
The propagator stems from $tg(\pi\alpha(t)/2).$ There are no poles (physical states) in the cross-channel of amplitudes of the ”alien” (in this case positive, with even $J$) signature.
This short prelude of long-known facts will be useful to fix designations in presenting our arguments further.
Odderon in the ”flat” implementation.
Problems in theory and the data processing.
Theory
At ”high enough” energies it is usually supposed that the contributions of secondary trajectories are negligibly small. In this case, it is assumed, as made also in the paper [2], that only the vacuum trajectory, Pomeron, $\alpha_{\mathcal{P}}(t)$, and its $C$-odd partner, Odderon, $\alpha(t)$ remain. A thorough analysis of the available data (from the ISR to the LHC) on $pp$ and $\bar{p}p$ interactions was undertaken in Ref.[2] in order to assess the scale of the influence of the Odderon exchange against the background of the Pomeron dominance.
In relation to the Odderon the authors, following in part Ref.[4], took an extreme position, conjecturing a flat Odderon ”trajectory”
$\alpha(t)=1.$ (3)
Direct use of ”trajectory” (3) in Eq.(1) is not possible so we proceed with use of a ”regularization”
$\alpha(t)\rightarrow 1+\alpha^{{}^{\prime}}t$ (4)
at low $t$ and then monitor the consequences. Taking afterwards $\alpha^{{}^{\prime}}=0$ we reproduce the ”flat Odderon” premise of Ref.[2]. Substiting $\alpha(t)\rightarrow 1+\alpha^{{}^{\prime}}t$ in Eq.(1) and considering it in the limit $\alpha^{{}^{\prime}}\rightarrow 0,t\rightarrow 0$ we get
$\mathcal{O}(s,t)\sim\frac{\Gamma_{1}^{2}(t)P_{1}(z_{t})}{-t}$ (5)
Equation (5) means (cf.Eq.(2)) nothing more than the presence of a massless vector hadron in the $p$-wave partial amplitude of the $t$-channel since in this case the amplitude of the negative signature is physical (odd angular momentum, $J=1$).
As far as is known, such hadrons have not been observed and their appearance can be avoided only by making the the vertex $\Gamma_{1}^{2}(t)$ vanishing at $t=0$. Moreover, due to analyticity of $\Gamma_{1}(t)$, zeroing can only be of the type $\Gamma_{1}^{2}(t)\rightarrow$ const $t^{2N}$ where $N\geq 1$ is an integer .
So we see that at the level of the Born amplitude the Odderon decouples from the proton at $t=0$. The real part decouples no weaker than tsimilar-$\sim t$, while the imaginary part decouples no weaker than $\sim t^{2}$.
Of course, what was said above about the massless pole applies only to the Born term, but constructing the full eikonal series requires integrating the products of the Born terms over the momentum transfers $t_{i}$ including $t_{i}=0$ , which leads to logarithmic divergences. The trick with $\Gamma_{1}^{2}(t)\sim t^{2}$ could ”save” the situation, but further discussion of this option lies beyond our prerogatives.
The data description in the ”flat” Odderon scenario
Leaving aside criticism of the purely theoretical foundations of the Odderon part of the paper [2], let us nevertheless consider its descriptive part in relation to the available data on elastic pp scattering.
It is important to note that authors provide both the central values of the fitted parameters of the models and their errors. However, the parameters extracted from the models (for example, the parameter $\rho$ [6] ) are provided only with their central values; for some reason, the authors do not provide the errors which can be of a paramount importance for the final conclusions. Namely, in the experimental data at the $t$ intervals the values of $\mid t\mid$ were carefully selected based on the high confidence level (at least $80\%$) for the extracted parameters. In particular, the work [5] shows that $\rho=0.01\pm 0.04$, i.e. the extracted value is determined with an accuracy of $40\%$(p. 054003-9). This is fundamentally at odds with the error of the authors of the TOTEM experiment, who groundlessly claim an error in the $\rho$parameter to be only $10\%$(with the same central value). It is clear that with such an error value, it is at least premature to talk about a discrepancy with the predictions of the COMPETE collaboration[7]. So, arguing about the difference in the values of the parameters in this work is generally pointless without indication of the errors. The authors of [2] refer to the work just mentioned [5] , but they cite intermediate values of the parameter $\rho$ from this work, and not its final value and its error from the conclusions of the referenced article. Moreover, the authors, referring to work[5], fairly note that its results were obtained on the basis of the TOTEM experiment only. Just because the official results of the ATLAS experiment did not yet exist at that time.
As soon as the latter appeared, they were studied with the same care as in work [5]. These results are presented, for example, in Ref. [8] , to which the authors of [2] also refer. The conclusions of this work show that the extracted parameters $\sigma_{tot}$ and $\rho$ have significantly lower values than in the TOTEM experiment. Moreover, the parameter $\rho$ itself changes significantly depending on the working array of experimental parameters used to extract it (always at a confidence level significantly higher than $50\%$) and thus even in a single ATLAS experiment it is extracted significantly ambiguously. In any case, it is essentially lower than that in the TOTEM experiment.
Thus, these two experiments contradict each other on the essence. Either one of them is incorrect, or they are both incorrect.
The inconsistency of the differential cross-section measurements in these experiments is already evident from the sets of their experimental data. The distance between the central points of these experiments (at different values of $t$) can reach $2.5$ full standard deviations of the TOTEM and more than $8$ full standard deviations of the ATLAS. In either case, these are unacceptably large values to present the same physical quantity.
The ambiguity of measurements of differential cross sections in two different experimental facilities is also observed at $\sqrt{s}=7$ and $8$ TeV, which means that the extracted values would differ at these energies. Moreover, this ambiguity is most pronounced at $\sqrt{s}=13$ TeV. The results of measurements of the same physical quantity at the LHC at the same energies in two different experiments show that even at lower energies, where there were no two devices measuring differential cross sections, their values could differ from the true values. For this reason, all joint fits at different energies should always be carried out, for example, using some factors for the experimental data, their own for each experiment, and which are considered as fit parameters. This would allow artificially reducing very different experimental data to a more or less uniform array.
In fact, the authors of this work do exactly that, although they attribute the fitted multipliers not to experimental data, but to a theoretical function for differential cross-sections. At the same time, they did not bother to explain how this operation will (or will not) be reflected in the extracted parameters (and their errors). It seems to us that it would be correct to be concerned with the accuracy and correctness of the very experimental measurements, rather than engage in arithmetic exercises in describing unacceptably different measurements of the same physical parameters, and therefore incorrectly representing them.
Another remark to the description of the results of this work. Its authors provide the extracted parameters, adding to them data on the number of degrees of freedom and the value of the total $\chi^{2}/pdf$.
This allows us to calculate the confidence level of this extraction of parameters. So, in more than half of the cases, this confidence level is extremely low or significantly lower than $50\%$. On such a background claims that ”the presence of the C-odd(Odderon)contribution essentially improves” the quality of the data description seem speculative and unfounded.
Conclusions
Summarizing all the above said, we have to conclude that, unfortunately, both the theoretical basis and the statistical arguments in the paper [2] cannot be considered sound. Accordingly, the same goes for their physical findings.
We hope that our remarks could be helpful for improving the further studies in this field.